OpenStudy (anonymous):

Check? :) If 5b − 25 = 25, what is the value of b? A. There is no solution. B. 0 C. 10 <--- D. 50

hartnn (hartnn):

thats correct :)

OpenStudy (imtiaz7):

5b=25+25 5b=50 b=50/5 b=10

OpenStudy (anonymous):

50-25=25 50=5*10